OpenStudy (anonymous):

how do you solve for -17 mod 7

OpenStudy (anonymous):

-17 = -17 + 21 = 4.

OpenStudy (anonymous):

what is the rule? why is it not -17 + 14 = -3?

OpenStudy (anonymous):

That's also true. I assumed you wanted an answer that was an element of {0,1,2,3,4,5,6}